a) a farmer sells cabbages for N$ 10 per head. The farmers variable costs are N$ 2.50 per head and total cost of 100 heads is N$1450.

I) how many cabbages must the manufacturer produce each month to break even?

II) how many cabbages should be produced to make profit

b) Total cost of producing carrot is C(X)=3600+100x+2x^2 and the total revenue function R(x)= 500x-2x^2

I) find the number of kg that maximizes profit

II) find maximum profit







1
Expert's answer
2021-10-01T16:54:20-0400

a)

i

Given

Price = 10

Variable cost = 2.50

Total cost of 100=1450



Break even point

"TC=TR\\

1450=10Q\\

Q=145"



ii


We know that

"TC=VC+FC\\

1450=(2.50\times 100)+FC\\

1450=250+FC\\

FC=1200"


Inorder to make profit 161 cabbages must be produced.


"P=TR-TC\\

P=(10\times 161)-(1200+(161\times 2.5))=7.5"


b)

i

"TR=TC\\500x-2x^2=3600+100x+2x^2\\4x^2-400x+3600=0\\4
(
𝑥
^2
−
1
0
0
𝑥
+
9
0
0
)
=
0\\x=10\\x=90"


90kg

ii

profit

"=price\times quantity\\90\times 7.5=675"